OpenStudy (anonymous):

A number decreased by -3 is equal to -40. Write an equation.

OpenStudy (hihi67):

x-3=-40 +40 x+37=0 -37 x=-37

OpenStudy (anonymous):

A. -3 - n = - 40 B. n - 40 = -3 C. n - (-3) = -40 D. n + (-3) = - 40

OpenStudy (hihi67):

D

OpenStudy (hihi67):

-37+(-3)=-37-3=-40